### v1

# A non-Hermitian loop for a quantum measurement

Abstract
We establish a framework for a mechanism steering state vector collapse through time evolution as described by the Schrödinger’s equation. Under the premise of a non-Hermitian Hamiltonian dynamics, a main outcome is that collapse has to occur when the Hamiltonian completes a closed loop in the parameter space encoding the interaction with the meter. For two-level systems, we put forward the phenomenon of chiral state conversion as a mechanism effectively eliminating superpositions. Our results open a way to simulate quantum measurements in classical systems that up to now were restricted to the Schrödinger part of the quantum dynamics.
Introduction and overview of key findings.– One century after its inception, quantum mechanics stands as one of humanity’s most rigorously tested theories, yet it still keeps a foundational challenge known as the measurement problem Bell (1990); Dirac (1963); Laloë (2001). The crux of the measurement problem lies in explaining or making sense of the wave function collapse. The postulates of quantum mechanics state that during a measurement, the Schrödinger equation, which ordinarily governs quantum dynamics, is momentarily suspended to allow the wave function to collapse (see for example Cohen-Tannoudji et al. (2020); Shankar (1994)). This split description of the quantum dynamics is introduced to bridge the complex world of quantum superpositions with the familiar classical reality, where those superpositions have vanished. The incompatibility between the two descriptions for the dynamics entails a “cut” between the quantum and classical realms (as Werner Heisenberg called it in a posthumously published letter Bacciagaluppi and Crull (2011) in response to the EPR paper Einstein et al. (1935)). Indeed, the collapse is essentially a non-unitary, irreversible operation with loss of information and eventual energy and momentum exchange Pearle (2000), or as John Wheeler put it “an irreversible act of amplification such as the blackening of a grain of silver bromide emulsion” Wheeler (2016)  [^1] Wheeler mentions that the term irreversible amplification was used by Bohr, see chapter 1 of Bohr (2010), which was published in 1958..
Despite advances in understanding the quantum-classical transition Zurek (2003); Schlosshauer (2007), the measurement problem remains a mystery for many of us Leggett (2006); Hance and Hossenfelder (2022)  [^2] There are also notable opinions against thinking of the measurement problem as such, see for example Mermin (2022). We also refer to Refs. Schlosshauer (2011); Schlosshauer et al. (2013) for a broader overview on the different positions.. Two proposals to address the conventional view of measurements are the Ghirardi-Rimini-Weber (GRW) Ghirardi et al. (1986, 1990) and the Continuous Spontaneous Localization (CSL) models Pearle (1989). Both introduce stochastic elements to the Schrödinger’s equation, suggesting that wave function collapse can occur spontaneously Bassi and Ghirardi (2003) (there are also other proposals using stochastic elements such as Fernández-Alcázar and Pastawski (2015)). Although GRW models do not rely on the concept of an apparatus to induce collapse, they still lack a microscopic explanation for how collapse occurs. Other proposals, inspired by GRW, are due to Diosi Diósi (1989); Donadi et al. (2021) and Penrose Penrose (1996) linking the wave function collapse to gravity. More recently, different authors (see  Martínez Romeral et al. (2024); Singh et al. (2023a, b) and also Van Wezel (2010); Mertens et al. (2023); Mukherjee and van Wezel (2024)) proposed using a non-Hermitian Hamiltonian to model a quantum measurement (earlier proposals also include dissipative modifications of the Schrödinger equation Gisin and Piron (1981)). These models are steps to make sense of quantum collapse Bell (1990), yet they also underscore its complexity Bassi et al. (2013) and the need for further experiments.
In a parallel and unfolding story, over recent years, we have seen a rise in research fronts using photonics Rechtsman et al. (2013); Vicencio et al. (2015); Guzmán-Silva et al. (2014); Stützer et al. (2018), metamaterials Veenstra et al. (2024), acoustics Zhang et al. (2021), mechanical systems Lorenz et al. (2023) and circuits Lee et al. (2018); Imhof et al. (2018) as classical analogs for quantum systems. Notably, in photonic lattices, the paraxial Helmholtz equation for the electromagnetic waves maps to the Schrödinger equation Szameit and Nolte (2010) thus allowing to simulate quantum systems. Similarly, in electrical circuits, the correspondence between the circuit Laplacian and the tight-binding Hamiltonian have been extensively used to investigate topological states Lee et al. (2018); Imhof et al. (2018). Moreover, it has been demonstrated that these circuits can also mimic Schrödinger dynamics Ezawa (2019). Collectively, these examples illustrate the availability of a classical toolkit for quantum system simulation. However, a noteworthy limitation remains apparent: classical simulators have yet to fully address quantum measurements, which, because of the split description mentioned earlier, are distinct from quantum dynamics and thus remain a significant challenge.
In this Letter, we present a theoretical framework for non-Hermitian models of wave function collapse. Non-Hermitian Hamiltonians have been shown to exhibit many interesting features Moiseyev (2011); Longhi (2009); Rotter and Bird (2015). One of them is the appearance of non-Hermitian degeneracies, called exceptional points Kato (1995); Moiseyev (2011), where both eigenvalues and eigenvectors coalesce. Exceptional points are at the heart of many counterintuitive Heiss and Harney (2001); Dembowski et al. (2001); Heiss (2012) and useful phenomena, such as the typical amplification of small parametric variations found near them which may lead to novel sensors  Wiersig (2014); Rechtsman (2017); Hodaei et al. (2017); Chen et al. (2017); Hokmabadi et al. (2019). Furthermore, non-Hermitian Hamiltonians offer a multitude of testing grounds in metamaterials Veenstra et al. (2024), photonics Szameit et al. (2011); Regensburger et al. (2012); Pan et al. (2018); Miri and Alù (2019); Vicencio et al. (2024), acoustics Zhang et al. (2021) and circuits Schindler et al. (2011); Hofmann et al. (2020), thereby offering a unique opportunity of simulating and discussing quantum collapse outside of the purely quantum domain.
To set the stage, we first revisit the postulates of quantum mechanics, laying the groundwork for our discussion. Subsequently, we discuss a set of physically motivated assumptions and their consequences. The main outcome of this part is that, if a quantum measurement is to be modeled as a process that occurs in time and which is described by the Schrödinger equation with a non-Hermitian Hamiltonian encoding the effect of the meter (or whatever interaction leading to collapse), then the Hamiltonian should describe a closed loop in parameter space. Interestingly, we are able to readily link this discovery with the phenomenon called state exchange or chiral state conversion: For the $2\times 2$ case, when moving along a closed contour in the space of the Hamiltonian parameters, the initial state undergoes a conversion to a preferred state, irrespective of initial conditions, which is determined by the sense of rotation in the space of parameters Uzdin et al. (2011); Berry and Uzdin (2011); Hassan et al. (2017a). Thus, chiral state conversion serves the purpose of breaking the superpositions and modeling quantum collapse under a single dynamic equation. However, there is a twist to consider: the sense of the loop is a property of the interaction itself, rather than the initial Hermitian Hamiltonian. In other words, it is an additional variable, and compatibility with standard quantum mechanics requires that its value must be set to satisfy Born’s rule. We close by discussing the implications of our results in the path to simulating a quantum measurement in classical systems.
Split procedures for the quantum dynamics.– Quantum mechanics is traditionally based on four key postulates, as outlined in textbooks Cohen-Tannoudji et al. (2020); Shankar (1994). Two of these postulates specifically address the theory’s dynamics. One of them pertains to the standard, measurement-free behavior governed by the Schrödinger equation, highlighting how systems evolve when not being measured. The other postulate introduces the concept of measurement as a distinct phenomenon, where the act of measuring disrupts the usual dynamics, causing the system’s state vector to collapse into one of the observable’s eigenstates. This collapse results in the system adopting a state that corresponds to one of the observable’s eigenvalues, rendering the measurement as an exceptional process within quantum mechanics, described by a different antagonistic procedure.
Figure 1: (a) Scheme showing a loop in a multidimensional parameter space defining the Hamiltonian. ${\cal H}_{0}$ is the initial Hamiltonian just before the interaction that leads to the collapse starts. (b) Time-evolution of the spin-up (blue) and spin-down (orange) relative probabilities for a system evolving under a Hamiltonian proportional to $\sigma_{x}$. A non-Hermitian term adding gains on the spin-up and loss on the spin-down is turned on with a gaussian function centered at time $t=5$ (standard deviation is $1$). This pulsed perturbation leads to a collapse after which the time-evolution restarts. (c) Scheme showing the Riemann surfaces of the real and imaginary part of the eigenvalues near an exceptional point. (d) Because of the peculiar shape of the solutions in presence of an exceptional point, the phenomenon of chiral state conversion leads to the asymmetric conversion of the initial eigenstates |$\ket{\psi_{1}}$ and |$\ket{\psi_{2}}$ after a closed loop which defies the adiabatic intuition and destroys superposition states.
Assumptions for a dynamical model of a quantum measurement.– Our discussion is based on the following set of assumptions for a family of models based on non-Hermitian Hamiltonians:
  1. 1.
    The measurement takes place over a finite time and the time evolution is dictated by the Schrödinger equation, with an eventual renormalization to keep the norm during the measurement process.
  2. 2.
    Over the timescale of the measurement, the Hamiltonian describes a closed loop in parameter space $\mathbf{X}(t)=\{X_{1}(t),X_{2}(t),...\}$, so that it remains unchanged at the end of the process (Fig. 1 (a)):
    $\begin{split}{\cal H}(\mathbf{X}(t=T))={\cal H}(\mathbf{X}(t=0))={\cal H}_{0}.\end{split}$ (1)
  3. 3.
    The interaction between the system and the measuring apparatus is modeled by introducing a non-Hermitian term to the Hamiltonian. For consistency with the previous points, this term needs to be time-dependent and describe a loop in parameter space. Details of the loop may also depend on the initial state.
Assumption 1 is the most natural one in order to recover unity in the description of the dynamics. Assumption 2 seems a reasonable assumption as we want for the system to remain the same when the apparatus (or the source of interaction) is removed. As a byproduct, the idea of making a closed loop in parameter space is naturally introduced.
Regarding assumption 3, the non-Hermitian term is a minimal model to account for the action of the measurement apparatus on the system without recurring to a larger Hilbert space, probably containing the degrees of freedom of the apparatus and/or the more complicated many-body interactions that escape the simple single-particle picture Foa Torres (2019). Another, more pragmatic motivation is that non-Hermitian Hamiltonians are readily available in quantum simulators using photonic lattices Szameit et al. (2011); Regensburger et al. (2012); Pan et al. (2018); Miri and Alù (2019) and electrical circuits Schindler et al. (2011); Hofmann et al. (2020).
All these assumptions can be summarized in two sentences: The measurement takes place over a finite time and is modeled as a non-Hermitian perturbation to an otherwise Hermitian Hamiltonian ${\mathcal{H}}_{0}$ which describes a loop in parameter space, thereby returning to the initial value after the interaction is complete. The parameter space is composed of those variables defining the Hamiltonian (including those related to the meter) and some of these parameters may have values set by the initial state.
The case of Ref. Martínez Romeral et al. (2024), where the collapse is obtained in a Stern-Gerlach setup (including spin and position variables), does not include an explicit time-dependence. However it can be thought as a particular equivalent case since its role is taken by the passage of the particle through the region where the magnetic field is acting.
Possible mechanisms.– Let us now turn to the discussion of the possible mechanisms. There are different ways where a non-Hermitian term could lead to a collapse of a superposition state on one of its components. A very simple way would be just adding gains and losses as to amplify one of the possible states and reduce the others. A typical case is shown in Fig.1 (b), where a superposition state is collapsed onto one of its components. In this case, we take for illustration purposes:
${\cal H}={\cal H}_{0}+\lambda{\cal H}_{1}.$ (2)
where:
${\cal H}_{0}=\gamma_{0}\begin{pmatrix}0&1\\ 1&0\end{pmatrix},$ (3)
${\cal H}_{1}=\begin{pmatrix}i\Gamma&0\\ 0&-i\Gamma\end{pmatrix}.$ (4)
The time dependence is embedded in the parameter $\lambda$ which is taken as a normalized Gaussian centered at the measurement time ($t=5$ in units of $\hbar/\gamma_{0}$) and with a standard deviation $\sigma=1$. Fig.1 (b) shows the probabilities of having spin up or down (dashed and full lines respectively) as a function of time. The meter acts around $t=5$ amplifying the spin down and reducing the spin up components leading to a situation where the superposition vanishes, the collapse. After the interaction with the meter vanishes the usual time-evolution leads again to a superposition state.
While this scheme may be adjusted to obtain a desired outcome it leaves, however, no much insight on the process of the measurement itself nor the ultimate mechanism leading to the collapse. One may also argue that substantial fine tuning of the non-Hermitian Hamiltonian is necessary to mimic a quantum collapse  [^3] Note that in this case the Hamiltonian describes a loop in parameter space enclosing a vanishing area..
For simplicity, in what follows we will focus on energy measurements A hint for the search for other possible mechanisms, is that the underlying phenomenon needs to bear an intrinsic asymmetry so that by changing a variable in the “measurement apparatus” one could obtain the different possible outcomes, it also needs to be a destroyer of superpositions. Furthermore, the mathematical substrate for such phenomenon is likely to break many of the usual rules to serve as a proper backdrop for the cut. Within the realm of non-Hermitian systems, one of the most prominent sources of unusual and counter-intuitive physics are the so called exceptional points Kato (1995); Heiss (2012). These are points in parameter space where the eigenvalues and eigenvectors coalesce (see the scheme in Fig.1 (c) illustrating the Riemann surfaces close to an exceptional point), and have been the focus of intense research  Dembowski et al. (2001, 2004); Stehmann et al. (2004); Hu et al. (2017); Miri and Alù (2019). The coalescence at an exceptional point needs to be distinguished from a usual Hermitian degeneracy where a full set of eigenstates is kept. The key difference being the appearance of defectiveness Foa Torres (2019), the lack of a full set of eigenstates. The dependence of the eigenvalues on the parameters around exceptional points is typically non-analytic Heiss (2012).
Although for a $2\times 2$ system non-Hermitian system, the adiabatic solutions are expected to exchange when encircling an exceptional point Heiss and Harney (2001); Dembowski et al. (2001); Heiss (2012), this conversion has been predicted to occur in an asymmetric way Uzdin et al. (2011); Berry and Uzdin (2011), the so called chiral state or mode conversion Hassan et al. (2017a) because of non-adiabatic transitions. This phenomenon manifests in a two-state system during the encirclement of an exceptional point within the parameter space. In the realm of Hermitian systems, the adiabatic theorem posits that if a system is initially prepared in one of its eigenstates and the Hamiltonian parameters are gradually altered along a closed trajectory, the system will revert to its initial state, albeit with an additional geometric phase Berry (1984). In contrast, in non-Hermitian systems, a clockwise encirclement around an exceptional point invariably results in the system settling into one of the eigenstates (deemed the preferred state). On the other hand, an anticlockwise journey consistently leads to the emergence of the second eigenstate, as depicted in Fig.1 (d). This intriguing behavior has been observed across various platforms, including waveguides Doppler et al. (2016), and optomechanical systems Xu et al. (2016). At a fundamental level, this state conversion is associated with the Stokes phenomenon of asymptotics Berry and Uzdin (2011) and the concept of stability loss delay Milburn et al. (2015).
Remarkably, the dynamics of the instantaneous eigenstates’ weights in non-Hermitian systems diverge significantly from those in Hermitian counterparts. The weights can exhibit extreme fluctuations Berry and Uzdin (2011), and transitions can occur even under slow dynamics, challenging the conventional understanding of adiabaticity Doppler et al. (2016); Berry and Uzdin (2011). Ref. Hassan et al. (2017a) has presented analytical and numerical insights on why a preferred state gets amplified in a dynamical setting and has shown that the state conversion is robust to deformations of the contour if the initial and final points are kept Hassan et al. (2017a). Recent findings further reveal that chiral state conversion can occur even without the need to encircle the exceptional point Hassan et al. (2017b); Nasari et al. (2022), thus broadening the phenomenon’s observed scope.
In our present context, chiral state conversion provides a promising mechanism for modeling the state vector collapse in a $2\times 2$ system. For this, the interaction with the meter should lead to exceptional points in parameter space. The loop starts (and ends) with the Hamiltonian being ${\cal H}_{0}$ which is Hermitian. Right at the beginning of the loop, a crucial decision must be made regarding the direction in which the loop will proceed. Switching between clockwise and anticlockwise loops changes the outcome or the preferred state. Upon the loop’s completion, when the interaction ceases and the Hamiltonian reverts to its original form ${\cal H}_{0}$, the initial superpositions that may have been in the initial state disappear, ushering in a state of certainty. This transition marks a pivotal moment in the dynamics of the system, where the ephemeral nature of quantum superpositions gives way to definitive outcomes, encapsulating the essence of quantum measurement and evolution.
The probability with which the meter chooses between the clockwise and anticlockwise loops should depend on the initial state and follow Born’s rule if the usual quantum mechanics predictions are to be preserved. This renders the dynamical equation not only non-Hermitian but also non-linear. Another point is that, for a sufficiently slow variation of the parameters, one would not need to fine tune the loop to encircle the exceptional point as the conversion has been shown to occur anyway Hassan et al. (2017b); Nasari et al. (2022).
Collapse when there is no measurement apparatus.– Collapse is known to occur even without a measurement apparatus. When there is a measurement apparatus it seems reasonable to model it as an additional non-Hermitian term that acts during the time of the measurement. But how to account for exceptional points in other cases, as when many body interactions are present? A possible path for an answer could be the proliferation of exceptional points reported in Luitz and Piazza (2019) or the phenomena reported in Žnidarič (2022).
We underline the fact that any interaction that vanishes after some time (this is, that does not modify permanently the Hamiltonian of the system of interest) adheres to assumption 2. This compliance allows for the autonomy of loops within the parameter space, enabling for spontaneous collapse, or collapse without a meter, to occur. In this case, one should just rather view collapse as emerging from the interaction with other system, encoded in an effective way in the loop in parameter space.
Classical simulators for state vector collapse.– As a corollary, and perhaps most immediately useful outcome of our discussion, one can envision the simulation of quantum measurements using classical objects: photonic setups with classical light, electrical circuits and many others. Indeed many existing experiments could be considered as first steps towards classical realizations/emulations of quantum measurements Dembowski et al. (2001, 2004); Doppler et al. (2016). The role of time could be played by a space coordinate along which the non-Hermitian term can be modulated as done in photonic systems Rechtsman et al. (2013).
Final remarks.– We have discussed a framework for a dynamical model of a quantum measurement where collapse emerges from an evolution dictated by the Schrödinger equation. A quantum measurement is seen as resulting from the interaction with other degrees of freedom introducing a non-Hermitian term to an otherwise Hermitian Hamiltonian. The nature of this interaction, having a finite lifespan, motivates using a time-dependent Hamiltonian which, throughout the interaction, describes a closed loop in parameter space. As a possible mechanism we signal the phenomenon called chiral state conversion, which is enabled by exceptional points and leads to the required destruction of superposition states. For a $2\times 2$ system, the outcome can be switched by changing the encircling of the loop from clockwise to anticlockwise. The probability of taking one option or the other, which amounts to choosing how the action of the meter takes place, needs to be given by the initial state so as to recover the usual predictions of quantum mechanics. This makes the dynamic equation also non-linear, which is consistent with other studies Mertens et al. (2021).
Chiral mode conversion has been extensively studied for $2\times 2$ systems, but there is still much to understand for systems with more degrees of freedom. Another mechanism that may come handy is the non-Hermitian skin effect Martinez Alvarez et al. (2018); Yao and Wang (2018), which essentially leads to an anomalous localization at one edge of a pristine lattice. In some cases one might tune the non-Hermitian term to select a few or even a single eigenstate. There has been a surge of interest in this phenomenon, which has been shown to have a topological origin Okuma et al. (2020); Zhang et al. (2020), see for example Okuma and Sato (2023); Bergholtz et al. (2021); Lin et al. (2023) and references therein.
We hope that our discussion may trigger a fresh look to this enduring enigma. Perhaps the challenge ahead is to take the measurement out of the quantum realm and down to an experiment with classical objects.
Acknowledgements.
L.E.F.F.T. thanks Joaquín Márquez Delgado, Diego Bautista Avilés, Igor Gornyi, Yuval Gefen, and Mikhail Kiselev for useful discussions, and C. Cormick for useful comments. L.E.F.F.T. acknowledges financial support by ANID FONDECYT (Chile) through grant 1211038, The Abdus Salam International Center for Theoretical Physics and the Simons Foundation, and by the EU Horizon 2020 research and innovation program under the Marie-Sklodowska-Curie Grant Agreement No. 873028 (HYDROTRONICS Project). ICN2 is funded by the CERCA programme / Generalitat de Catalunya, and is supported by the Severo Ochoa Centres of Excellence programme, Grant CEX2021-001214-S, funded by MCIN / AEI / 10.13039.501100011033. This work is also supported by MICIN with European funds‐NextGenerationEU (PRTR‐C17.I1) and by 2021 SGR 00997, funded by Generalitat de Catalunya.

### v2

@@
 Abstract
-We establish a framework for a mechanism steering state vector collapse through time evolution as described by the Schrödinger’s equation.
-Under the premise of a non-Hermitian Hamiltonian dynamics, a main outcome is that collapse has to occur when the Hamiltonian completes a closed loop in the parameter space encoding the interaction with the meter.
-For two-level systems, we put forward the phenomenon of chiral state conversion as a mechanism effectively eliminating superpositions.
-Our results open a way to simulate quantum measurements in classical systems that up to now were restricted to the Schrödinger part of the quantum dynamics.
-Introduction and overview of key findings.– One century after its inception, quantum mechanics stands as one of humanity’s most rigorously tested theories, yet it still keeps a foundational challenge known as the measurement problem Bell (1990); Dirac (1963); Laloë (2001).
+Here we present a non-Hermitian framework for modeling state-vector collapse under unified dynamics described by Schrödinger’s equation.
+Under the premise of non-Hermitian Hamiltonian dynamics, we argue that collapse has to occur when the Hamiltonian completes a closed loop in the parameter space encoding the interaction with the meter.
+For two-level systems, we put forward the phenomenon of chiral state conversion as a mechanism for effectively eliminating superpositions.
+This perspective opens a way to simulate quantum measurements in classical systems that up to now were restricted to the Schrödinger part of the quantum dynamics.
+I Introduction
+One century after its inception, quantum mechanics stands as one of humanity’s most rigorously tested theories, yet it still keeps a foundational challenge known as the measurement problem Bell (1990); Dirac (1963); Laloë (2001).
 The crux of the measurement problem lies in explaining or making sense of the wave function collapse.
-The postulates of quantum mechanics state that during a measurement, the Schrödinger equation, which ordinarily governs quantum dynamics, is momentarily suspended to allow the wave function to collapse (see for example Cohen-Tannoudji et al.
+The postulates of quantum mechanics state that during a measurement, the Schrödinger equation, which ordinarily governs quantum dynamics, is momentarily suspended to allow the wave function to collapse (see, for example, Cohen-Tannoudji et al.
 (2020); Shankar (1994)).
@@
 Despite advances in understanding the quantum-classical transition Zurek (2003); Schlosshauer (2007), the measurement problem remains a mystery for many of us Leggett (2006); Hance and Hossenfelder (2022)  [^2] There are also notable opinions against thinking of the measurement problem as such, see for example Mermin (2022).
-We also refer to Refs.
-Schlosshauer (2011); Schlosshauer et al.
-(2013) for a broader overview on the different positions..
+We also refer to Schlosshauer (2011) for a broader overview of the different positions.
 Two proposals to address the conventional view of measurements are the Ghirardi-Rimini-Weber (GRW) Ghirardi et al.
@@
 Although GRW models do not rely on the concept of an apparatus to induce collapse, they still lack a microscopic explanation for how collapse occurs.
-Other proposals, inspired by GRW, are due to Diosi Diósi (1989); Donadi et al.
+Other proposals, inspired by GRW, are due to Diósi Diósi (1989); Donadi et al.
 (2021) and Penrose Penrose (1996) linking the wave function collapse to gravity.
@@
 (2023); Mukherjee and van Wezel (2024)) proposed using a non-Hermitian Hamiltonian to model a quantum measurement (earlier proposals also include dissipative modifications of the Schrödinger equation Gisin and Piron (1981)).
-These models are steps to make sense of quantum collapse Bell (1990), yet they also underscore its complexity Bassi et al.
+These models are steps in making sense of quantum collapse Bell (1990), yet they also underscore its complexity Bassi et al.
 (2013) and the need for further experiments.
@@
 (2013); Vicencio et al.
-(2015); Guzmán-Silva et al.
-(2014); Stützer et al.
 (2015), metamaterials Veenstra et al.
@@
 (2023) and circuits Lee et al.
-(2018); Imhof et al.
 (2018) as classical analogs for quantum systems.
@@
 Similarly, in electrical circuits, the correspondence between the circuit Laplacian and the tight-binding Hamiltonian have been extensively used to investigate topological states Lee et al.
-(2018); Imhof et al.
 (2018).
 Moreover, it has been demonstrated that these circuits can also mimic Schrödinger dynamics Ezawa (2019).
-Collectively, these examples illustrate the availability of a classical toolkit for quantum system simulation.
-However, a noteworthy limitation remains apparent: classical simulators have yet to fully address quantum measurements, which, because of the split description mentioned earlier, are distinct from quantum dynamics and thus remain a significant challenge.
-In this Letter, we present a theoretical framework for non-Hermitian models of wave function collapse.
-Non-Hermitian Hamiltonians have been shown to exhibit many interesting features Moiseyev (2011); Longhi (2009); Rotter and Bird (2015).
-One of them is the appearance of non-Hermitian degeneracies, called exceptional points Kato (1995); Moiseyev (2011), where both eigenvalues and eigenvectors coalesce.
-Exceptional points are at the heart of many counterintuitive Heiss and Harney (2001); Dembowski et al.
-(2001); Heiss (2012) and useful phenomena, such as the typical amplification of small parametric variations found near them which may lead to novel sensors  Wiersig (2014); Rechtsman (2017); Hodaei et al.
+Collectively, these examples illustrate the availability of a classical toolkit for simulation of quantum systems.
+However, a critical limitation remains: while a few noteworthy exceptions offer solutions for specific cases Longhi (2006); Liu et al.
+(2023), classical simulators generally lack a broadly applicable counterpart to quantum measurements, thereby preventing the exploration of a wealth of phenomena such as measurement-enriched phases.
+The aforementioned limitation of classical simulators in handling quantum measurement shows the need for a deeper understanding of the measurement process itself, which might then inspire new simulation approaches.
+Focusing on recent proposals modeling collapse using non-Hermitian Hamiltonians, challenges remain.
+Although models like our prior work Martínez Romeral et al.
+(2024) show how non-Hermitian dynamics can amplify a chosen state, the need to construct the interaction specifically to favor a predetermined outcome perhaps offers less insight into the measurement process itself.
+Moreover, such approaches do not necessarily provide a dynamical explanation for why the possible outcomes of an energy measurement correspond specifically to the eigenstates of the system’s original Hamiltonian ($\mathcal{H}_{0}$).
+One might desire a mechanism where both the preference for eigenstates and the selection among them emerge dynamically from the system’s response to control or interaction.
+This motivates our exploration of an alternative framework centered on dynamically induced collapse via parameter loops around exceptional points, aiming to demonstrate how such dynamics can naturally establish the initial eigenstates as the unique stable outcomes and provide a topological mechanism for selecting one.
+The alternative framework we propose leverages key features of non-Hermitian Hamiltonians Moiseyev (2011); Bergholtz et al.
+(2021); Lin et al.
+(2023), systems whose unique properties offer a potential pathway to address the limitations discussed above.
+Non-Hermitian Hamiltonians exhibit phenomena absent in their Hermitian counterparts, most notably non-Hermitian degeneracies known as exceptional points (EPs) Kato (1995); Moiseyev (2011), where both eigenvalues and eigenvectors coalesce.
+These EPs are central to many counterintuitive phenomena Heiss and Harney (2001); Dembowski et al.
+(2001); Heiss (2012) and useful effects, including sensitivity enhancements relevant for novel sensors Wiersig (2014); Rechtsman (2017); Hodaei et al.
 (2017); Chen et al.
@@
 (2019).
-Furthermore, non-Hermitian Hamiltonians offer a multitude of testing grounds in metamaterials Veenstra et al.
+Crucially for our goal of dynamically explaining state selection, the behavior near EPs, particularly the state evolution during parameter loops encircling them Hassan et al.
+(2017a), forms the core of our proposed collapse mechanism.
+Furthermore, the physics of non-Hermitian systems and EPs can be readily explored in various experimental platforms, including metamaterials Veenstra et al.
 (2024), photonics Szameit et al.
-(2011); Regensburger et al.
 (2011); Pan et al.
@@
 (2025), acoustics Zhang et al.
-(2021) and circuits Schindler et al.
+(2021), and circuits Schindler et al.
 (2011); Hofmann et al.
-(2020), thereby offering a unique opportunity of simulating and discussing quantum collapse outside of the purely quantum domain.
+(2020).
+This experimental accessibility offers a unique opportunity not only to test the fundamental dynamics we propose but also potentially to simulate aspects of quantum collapse, as modeled here, within controllable classical systems.
 To set the stage, we first revisit the postulates of quantum mechanics, laying the groundwork for our discussion.
 Subsequently, we discuss a set of physically motivated assumptions and their consequences.
-The main outcome of this part is that, if a quantum measurement is to be modeled as a process that occurs in time and which is described by the Schrödinger equation with a non-Hermitian Hamiltonian encoding the effect of the meter (or whatever interaction leading to collapse), then the Hamiltonian should describe a closed loop in parameter space.
-Interestingly, we are able to readily link this discovery with the phenomenon called state exchange or chiral state conversion: For the $2\times 2$ case, when moving along a closed contour in the space of the Hamiltonian parameters, the initial state undergoes a conversion to a preferred state, irrespective of initial conditions, which is determined by the sense of rotation in the space of parameters Uzdin et al.
-(2011); Berry and Uzdin (2011); Hassan et al.
+The main outcome of this part is that if a quantum measurement is to be modeled as a process that occurs in time and which is described by the Schrödinger equation with a non-Hermitian Hamiltonian encoding the effect of the meter (or whatever interaction leading to collapse), then the Hamiltonian should describe a closed loop in parameter space (returning to its initial Hermitian form).
+Interestingly, we are able to readily link this discovery with the phenomenon called state exchange or chiral state conversion Hassan et al.
+(2017a): For the $2\times 2$ case, when moving along a closed contour in the space of the Hamiltonian parameters, and under certain conditions for the speed with which the loop is completed, the initial state undergoes a conversion to a preferred state, which is determined by the sense of rotation in the space of parameters Hassan et al.
 (2017a).
-Thus, chiral state conversion serves the purpose of breaking the superpositions and modeling quantum collapse under a single dynamic equation.
-However, there is a twist to consider: the sense of the loop is a property of the interaction itself, rather than the initial Hermitian Hamiltonian.
+Thus, chiral state conversion breaks the superpositions and models quantum collapse under a single dynamic equation.
+However, there is a twist to consider: the sense of the loop is a property of the interaction itself rather than the initial Hermitian Hamiltonian.
 In other words, it is an additional variable, and compatibility with standard quantum mechanics requires that its value must be set to satisfy Born’s rule.
-We close by discussing the implications of our results in the path to simulating a quantum measurement in classical systems.
-Split procedures for the quantum dynamics.– Quantum mechanics is traditionally based on four key postulates, as outlined in textbooks Cohen-Tannoudji et al.
+We close by discussing the implications of our results on the path to simulating quantum measurement in classical systems.
+II Split procedures for the quantum dynamics
+Quantum mechanics is traditionally based on four key postulates, as outlined in textbooks Cohen-Tannoudji et al.
 (2020); Shankar (1994).
@@
 This collapse results in the system adopting a state that corresponds to one of the observable’s eigenvalues, rendering the measurement as an exceptional process within quantum mechanics, described by a different antagonistic procedure.
+The measurement postulate in quantum mechanics, in its conventional form, can be broken down into three key questions: 1) What are the possible outcomes of a measurement?
+2) What happens to the system’s state after a measurement (state-vector collapse)?
+and 3) What are the probabilities of obtaining each outcome (Born’s rule)?
+Here, we primarily address the first and second question, offering a model for the dynamics of the collapse itself.
 Figure 1: (a) Scheme showing a loop in a multidimensional parameter space defining the Hamiltonian.
 ${\cal H}_{0}$ is the initial Hamiltonian just before the interaction that leads to the collapse starts.
-(b) Time-evolution of the spin-up (blue) and spin-down (orange) relative probabilities for a system evolving under a Hamiltonian proportional to $\sigma_{x}$.
+(b) Time-evolution of the spin-up (blue, continous) and spin-down (orange, dashed) relative probabilities for a system evolving under the Hamiltonian of Eq.
+2 with $\gamma_{0}=1$ and $\Gamma=50$.
+The state at $t=0$ corresponds to a particle with spin down.
 A non-Hermitian term adding gains on the spin-up and loss on the spin-down is turned on with a gaussian function centered at time $t=5$ (standard deviation is $1$).
@@
 (c) Scheme showing the Riemann surfaces of the real and imaginary part of the eigenvalues near an exceptional point.
-(d) Because of the peculiar shape of the solutions in presence of an exceptional point, the phenomenon of chiral state conversion leads to the asymmetric conversion of the initial eigenstates |$\ket{\psi_{1}}$ and |$\ket{\psi_{2}}$ after a closed loop which defies the adiabatic intuition and destroys superposition states.
-Assumptions for a dynamical model of a quantum measurement.– Our discussion is based on the following set of assumptions for a family of models based on non-Hermitian Hamiltonians:
+(d) Because of the peculiar structure of the solutions in presence of an exceptional point, the phenomenon of chiral state conversion leads to the asymmetric conversion of the initial eigenstates |$\ket{\psi_{1}}$ and |$\ket{\psi_{2}}$ after a closed loop which defies the adiabatic intuition and destroys superposition states.
+Dynamical Model of a Quantum Measurement.– Our model for a dynamical description of quantum measurement using non-Hermitian Hamiltonians rests on three key assumptions:
   1. 1.
-    The measurement takes place over a finite time and the time evolution is dictated by the Schrödinger equation, with an eventual renormalization to keep the norm during the measurement process.
+    Finite-Time Evolution: The measurement occurs over a finite time interval.
+The system’s state vector, denoted as $|\Psi(t)\rangle$, evolves according to the standard linear time-dependent Schrödinger equation:
+    $i\hbar\frac{\partial|\Psi(t)\rangle}{\partial t}=\mathcal{H}(t)|\Psi(t)\rangle,$
+    where $\mathcal{H}(t)$ is the (generally non-Hermitian) Hamiltonian encoding the system and its interaction with the measurement apparatus.
+The resulting state vector $|\Psi(t)\rangle$ is not necessarily normalized to unity; its components (i.e., projections onto a basis, such as $\langle\text{outcome}|\Psi(t)\rangle$) are interpreted as relative probability amplitudes.
+To obtain physical probabilities for specific outcomes at any given time $t$, the squared magnitudes of these relative probability amplitudes (yielding relative probabilities) are subsequently normalized by the total instantaneous norm $\langle\Psi(t)|\Psi(t)\rangle$ (i.e., $P_{\text{outcome}}(t)=|\langle\text{outcome}|\Psi(t)\rangle|^{2}/\langle\Psi(t)|\Psi(t)\rangle$).
+We note that alternative theoretical approaches exist where norm preservation is enforced continuously via non-linear modifications to the Schrödinger equation (see, e.g., Ref.
+Brody and Graefe (2012)).
   2. 2.
-    Over the timescale of the measurement, the Hamiltonian describes a closed loop in parameter space $\mathbf{X}(t)=\{X_{1}(t),X_{2}(t),...\}$, so that it remains unchanged at the end of the process (Fig.
-1 (a)):
-    $\begin{split}{\cal H}(\mathbf{X}(t=T))={\cal H}(\mathbf{X}(t=0))={\cal H}_{0}.\end{split}$ (1)
+    Closed Loop in Parameter Space: The Hamiltonian $\mathcal{H}(\mathbf{X}(t))$ traces a closed loop in parameter space $\mathbf{X}(t)=\{X_{1}(t),X_{2}(t),...\}$ during the measurement, ensuring that it returns to its initial form $\mathcal{H}_{0}$ by the end of the process (Fig.
+1(a)).
+This can be represented as:
+    $\mathcal{H}(\mathbf{X}(t=T))=\mathcal{H}(\mathbf{X}(t=0))=\mathcal{H}_{0}$ (1)
   3. 3.
-    The interaction between the system and the measuring apparatus is modeled by introducing a non-Hermitian term to the Hamiltonian.
-For consistency with the previous points, this term needs to be time-dependent and describe a loop in parameter space.
-Details of the loop may also depend on the initial state.
-Assumption 1 is the most natural one in order to recover unity in the description of the dynamics.
-Assumption 2 seems a reasonable assumption as we want for the system to remain the same when the apparatus (or the source of interaction) is removed.
-As a byproduct, the idea of making a closed loop in parameter space is naturally introduced.
-Regarding assumption 3, the non-Hermitian term is a minimal model to account for the action of the measurement apparatus on the system without recurring to a larger Hilbert space, probably containing the degrees of freedom of the apparatus and/or the more complicated many-body interactions that escape the simple single-particle picture Foa Torres (2019).
-Another, more pragmatic motivation is that non-Hermitian Hamiltonians are readily available in quantum simulators using photonic lattices Szameit et al.
-(2011); Regensburger et al.
+    Non-Hermitian term: The interaction with the measurement apparatus is modeled by a time-dependent non-Hermitian term in the Hamiltonian.
+This term also describes a loop in parameter space, potentially dependent on the initial state of the system.
+These assumptions ensure that the system returns to its original form after the measurement apparatus is removed.
+The non-Hermitian term serves as a minimal model for the apparatus’s influence, avoiding the need to explicitly include the apparatus’s degrees of freedom in the Hilbert space, which would lead to a more complex many-body problem, and is readily implemented in quantum simulators such as photonic lattices Szameit et al.
 (2011); Pan et al.
@@
 (2020).
-All these assumptions can be summarized in two sentences: The measurement takes place over a finite time and is modeled as a non-Hermitian perturbation to an otherwise Hermitian Hamiltonian ${\mathcal{H}}_{0}$ which describes a loop in parameter space, thereby returning to the initial value after the interaction is complete.
-The parameter space is composed of those variables defining the Hamiltonian (including those related to the meter) and some of these parameters may have values set by the initial state.
 The case of Ref.
 Martínez Romeral et al.
-(2024), where the collapse is obtained in a Stern-Gerlach setup (including spin and position variables), does not include an explicit time-dependence.
-However it can be thought as a particular equivalent case since its role is taken by the passage of the particle through the region where the magnetic field is acting.
-Possible mechanisms.– Let us now turn to the discussion of the possible mechanisms.
+(2024), where collapse is obtained in a Stern-Gerlach setup including both spin and position variables does not include an implicit change in time.
+However, it can be viewed as a particular case where the particle’s passage through the magnetic field region effectively introduces the time dependence when viewed on the spin variable.
+In essence, our model describes the measurement-induced collapse as resulting from a continuous (and not instantaneous) process, driven by a non-Hermitian perturbation on the initial Hermitian Hamiltonian, which could be mapped to a close loop dynamics in parameter space, dictated by the variables of the Hamiltonian, some of which may be influenced by the initial state.
+III Possible mechanisms
+Let us now turn to the discussion of the possible mechanisms.
 There are different ways where a non-Hermitian term could lead to a collapse of a superposition state on one of its components.
-A very simple way would be just adding gains and losses as to amplify one of the possible states and reduce the others.
+A very simple way would be just to add gains and losses to amplify a target state and reduce the others.
 A typical case is shown in Fig.1 (b), where a superposition state is collapsed onto one of its components.
@@
 ${\cal H}_{1}=\begin{pmatrix}i\Gamma&0\\ 0&-i\Gamma\end{pmatrix}.$ (4)
-The time dependence is embedded in the parameter $\lambda$ which is taken as a normalized Gaussian centered at the measurement time ($t=5$ in units of $\hbar/\gamma_{0}$) and with a standard deviation $\sigma=1$.
-Fig.1 (b) shows the probabilities of having spin up or down (dashed and full lines respectively) as a function of time.
-The meter acts around $t=5$ amplifying the spin down and reducing the spin up components leading to a situation where the superposition vanishes, the collapse.
+The time dependence is embedded in the parameter $\lambda$ which is taken as a normalized Gaussian centered at the measurement time ($t=5$ in units of $\hbar/\gamma_{0}$, represented with a vertical dashed line in Fig.1 (b)) and with a standard deviation $\sigma=1$.
+Specifically, we take $\lambda(t)=\exp(-(t-t_{0})^{2}/(2\sigma^{2}))/(\sigma\sqrt{2\pi})$.
+The state vector $|\Psi(t)\rangle$ evolves under $\mathcal{H}$ according to the linear Schrödinger equation as outlined in Section II.
+The probabilities displayed in Fig.1 (b) for having spin up (blue continuous line) and spin down (orange dashed line) as a function of time (the initial state is a spin down state) are then calculated from the components of $|\Psi(t)\rangle$ – whose projections onto the spin basis states are treated as relative probability amplitudes – by squaring their magnitudes and normalizing by the total instantaneous norm $\langle\Psi(t)|\Psi(t)\rangle$.
+The meter acts around $t=5$ amplifying the spin up and reducing the spin down components leading to a situation where the superposition vanishes, the collapse.
 After the interaction with the meter vanishes the usual time-evolution leads again to a superposition state.
-While this scheme may be adjusted to obtain a desired outcome it leaves, however, no much insight on the process of the measurement itself nor the ultimate mechanism leading to the collapse.
-One may also argue that substantial fine tuning of the non-Hermitian Hamiltonian is necessary to mimic a quantum collapse  [^3] Note that in this case the Hamiltonian describes a loop in parameter space enclosing a vanishing area..
-For simplicity, in what follows we will focus on energy measurements A hint for the search for other possible mechanisms, is that the underlying phenomenon needs to bear an intrinsic asymmetry so that by changing a variable in the “measurement apparatus” one could obtain the different possible outcomes, it also needs to be a destroyer of superpositions.
+Although these models show how non-Hermitian Hamiltonians can amplify a chosen state, the need to construct the interaction specifically to favor a predetermined outcome perhaps offers less insight into the measurement process itself.
+One might desire a mechanism where state selection emerges more dynamically from the system’s response to control or interaction.
+This motivates our exploration of an alternative framework centered on dynamically induced collapse.
+For simplicity, in what follows we will focus on energy measurements.
+A hint for the search for other possible mechanisms, is that the underlying phenomenon needs to bear an intrinsic asymmetry so that by changing a variable in the “measurement apparatus” one could obtain the different possible outcomes, it also needs to be a destroyer of superpositions.
 Furthermore, the mathematical substrate for such phenomenon is likely to break many of the usual rules to serve as a proper backdrop for the cut.
@@
 These are points in parameter space where the eigenvalues and eigenvectors coalesce (see the scheme in Fig.1 (c) illustrating the Riemann surfaces close to an exceptional point), and have been the focus of intense research  Dembowski et al.
-(2001, 2004); Stehmann et al.
-(2004); Hu et al.
-(2017); Miri and Alù (2019).
+(2001, 2004); Miri and Alù (2019).
 The coalescence at an exceptional point needs to be distinguished from a usual Hermitian degeneracy where a full set of eigenstates is kept.
@@
 (2016).
-At a fundamental level, this state conversion is associated with the Stokes phenomenon of asymptotics Berry and Uzdin (2011) and the concept of stability loss delay Milburn et al.
+At a fundamental level, this state conversion has been associated with the Stokes phenomenon of asymptotics Berry and Uzdin (2011) and the concept of stability loss delay Milburn et al.
 (2015).
@@
 Hassan et al.
-(2017a) has presented analytical and numerical insights on why a preferred state gets amplified in a dynamical setting and has shown that the state conversion is robust to deformations of the contour if the initial and final points are kept Hassan et al.
-(2017a).
-Recent findings further reveal that chiral state conversion can occur even without the need to encircle the exceptional point Hassan et al.
+(2017a) has presented analytical and numerical insights on why a preferred state gets amplified in a dynamical setting.
+Recent papers indicate that chiral state conversion can occur even without the need to encircle the exceptional point Hassan et al.
 (2017b); Nasari et al.
 (2022), thus broadening the phenomenon’s observed scope.
-In our present context, chiral state conversion provides a promising mechanism for modeling the state vector collapse in a $2\times 2$ system.
+But the understanding of the phenomenon of state conversion is still far from complete, as recent studies indicate Nye (2023); Nye and Kantartzis (2024); Kumar et al.
+(2025).
+Clearly, more research is needed to get a better picture of the time scales needed for it to happen and its robustness.
+Chiral state conversion provides a promising mechanism for modeling the state vector collapse in a $2\times 2$ system.
 For this, the interaction with the meter should lead to exceptional points in parameter space.
@@
 This transition marks a pivotal moment in the dynamics of the system, where the ephemeral nature of quantum superpositions gives way to definitive outcomes, encapsulating the essence of quantum measurement and evolution.
-The probability with which the meter chooses between the clockwise and anticlockwise loops should depend on the initial state and follow Born’s rule if the usual quantum mechanics predictions are to be preserved.
-This renders the dynamical equation not only non-Hermitian but also non-linear.
-Another point is that, for a sufficiently slow variation of the parameters, one would not need to fine tune the loop to encircle the exceptional point as the conversion has been shown to occur anyway Hassan et al.
+To make the concept of forming a closed loop in parameter space more concrete, let us consider a minimal example which captures the essential physics of EP encirclement.
+A general form for such a Hamiltonian can be written relative to its parameters $\delta(t)$ and $\gamma(t)$ as:
+${\cal H}(t)=\begin{pmatrix}\delta(t)+i\gamma(t)&J\\ J&-\delta(t)-i\gamma(t)\end{pmatrix}$ (5)
+Here, $\mathbb{I}$ is the identity matrix and $J$ represents a real coupling strength, and $\delta(t)$ is a time-varying detuning.
+The term $\gamma(t)$ represents half the time-varying gain/loss difference between the two states.
+Its significance is central to the non-Hermitian nature of the Hamiltonian during the measurement interaction; it explicitly introduces gain or loss, modeling the strength and character of the system-apparatus coupling that drives the non-unitary evolution.
+The specific choice and time-dependence of $\gamma(t)$ (along with $\delta(t)$) define the parametric loop in the non-Hermitian parameter space, which is crucial for the proposed chiral state conversion mechanism.
+The model requires $\gamma(t)$ to be non-zero only during the transient measurement interaction, returning to zero (as does $\delta(t)$) at the beginning ($t=0$) and end ($t=T$) of the loop, ensuring the Hamiltonian reverts to the Hermitian $\mathcal{H}_{0}$.
+Let the initial and final Hamiltonian be the Hermitian ${\cal H}_{0}$, corresponding to parameters $\delta=0$, and $\gamma=0$:
+${\cal H}_{0}=\begin{pmatrix}0&J\\ J&0\end{pmatrix}.$ (6)
+The EP occurs at $(\delta_{EP},\gamma_{EP})=(0,J)$ (assuming $J>0$).
+A closed loop in the $(\delta,\gamma)$ parameter space that starts from ${\cal H}_{0}$ (i.e., starts at $(\delta=0,\gamma=0)$), encircles this EP, and returns to ${\cal H}_{0}$ at time $t=T$ can be explicitly implemented.
+For example, consider a circular path centered precisely at the EP $(0,J)$ with radius $R=J$.
+This path passes through the origin and is parameterized during the interval $0\leq t\leq T$ by:
+$\delta(t)=J\sin(\Omega t)$ (7)
+and
+$\gamma(t)=J(1-\cos(\Omega t)),$ (8)
+where $\Omega=2\pi/T$ determines the loop duration $T$.
+Note that at $t=0$ and $t=T$, we have $\delta=0$ and $\gamma=0$, ensuring the Hamiltonian starts and ends at ${\cal H}_{0}$.
+The time evolution governed by this explicit ${\cal H}(t)$ generates the phenomenon of chiral state conversion Hassan et al.
+(2017a).
+[…]
+To connect our model to the probabilistic outcomes of quantum measurements, we must address how the choice of loop direction (chirality) relates to Born’s rule.
+Our framework proposes that the non-Hermitian interaction drives the system towards a specific eigenstate depending on the chirality (loop direction).
+However, the model in its current form does not inherently derive the probabilities associated with these outcomes.
+To recover the predictions of conventional quantum mechanics, we must impose Born’s rule externally: the probability of the interaction selecting a particular chirality (clockwise or anti-clockwise) must be explicitly set to match the squared amplitude of the corresponding component in the system’s state vector just before the measurement interaction begins.
+The precise physical mechanism responsible for enforcing this probabilistic choice based on the initial state remains an open question not solved by this dynamical model.
+Requiring the chirality selection probability to depend on the initial state amplitudes inherently introduces a non-linear element into the overall measurement process description, distinguishing it from the purely linear Schrödinger evolution between measurements.
+This aligns with arguments suggesting that a purely linear dynamics cannot fully replicate Born’s rule  Mertens et al.
+(2021).
+Importantly, for a sufficiently slow variation of the parameters, encircling the exceptional point might not be necessary, as the conversion has been signaled to occur anyway Hassan et al.
 (2017b); Nasari et al.
 (2022).
-Collapse when there is no measurement apparatus.– Collapse is known to occur even without a measurement apparatus.
-When there is a measurement apparatus it seems reasonable to model it as an additional non-Hermitian term that acts during the time of the measurement.
-But how to account for exceptional points in other cases, as when many body interactions are present?
-A possible path for an answer could be the proliferation of exceptional points reported in Luitz and Piazza (2019) or the phenomena reported in Žnidarič (2022).
-We underline the fact that any interaction that vanishes after some time (this is, that does not modify permanently the Hamiltonian of the system of interest) adheres to assumption 2.
-This compliance allows for the autonomy of loops within the parameter space, enabling for spontaneous collapse, or collapse without a meter, to occur.
-In this case, one should just rather view collapse as emerging from the interaction with other system, encoded in an effective way in the loop in parameter space.
-Classical simulators for state vector collapse.– As a corollary, and perhaps most immediately useful outcome of our discussion, one can envision the simulation of quantum measurements using classical objects: photonic setups with classical light, electrical circuits and many others.
-Indeed many existing experiments could be considered as first steps towards classical realizations/emulations of quantum measurements Dembowski et al.
-(2001, 2004); Doppler et al.
-(2016).
-The role of time could be played by a space coordinate along which the non-Hermitian term can be modulated as done in photonic systems Rechtsman et al.
-(2013).
-Final remarks.– We have discussed a framework for a dynamical model of a quantum measurement where collapse emerges from an evolution dictated by the Schrödinger equation.
+This robustness may simplify the experimental implementation of our proposed mechanism.
+Looking ahead, a crucial question is how the probabilistic nature of quantum measurement emerges from the seemingly deterministic chiral dynamics.
+We speculate that, as long as the environmental interactions (at the origin of the non-Hermitian term) are random and unbiased, they will lead to a statistical distribution of chirality choices that mirrors the probabilities encoded in the quantum state, hence statistically aligning with Born’s rule and the predictions of conventional quantum mechanics.
+Exploring this connection between environmental interactions, non-Hermitian dynamics, and the emergence of Born’s rule is a promising direction for future research.
+Collapse when there is no measurement apparatus.– The collapse of the wave function can occur even in the absence of a dedicated measurement apparatus.
+While our model focuses on the case where a non-Hermitian term represents the measurement interaction, extending it to other scenarios, such as those involving many-body interactions, remains an open question.
+A possible path forward could involve exploring the proliferation of exceptional points Luitz and Piazza (2019) or the phenomena reported in Žnidarič (2022).
+Importantly, any interaction that vanishes after a finite time, leaving the system’s Hamiltonian unchanged, automatically satisfies our assumption 2, allowing for the possibility of spontaneous collapse driven by transient interactions represented as loops in parameter space.
+IV Final remarks
+We have discussed a framework for a dynamical model of a quantum measurement where collapse emerges from an evolution dictated by the Schrödinger equation.
 A quantum measurement is seen as resulting from the interaction with other degrees of freedom introducing a non-Hermitian term to an otherwise Hermitian Hamiltonian.
 The nature of this interaction, having a finite lifespan, motivates using a time-dependent Hamiltonian which, throughout the interaction, describes a closed loop in parameter space.
-As a possible mechanism we signal the phenomenon called chiral state conversion, which is enabled by exceptional points and leads to the required destruction of superposition states.
+As a possible mechanism, we signal the phenomenon called chiral state conversion, which leads to the destruction of superposition states in favor of one of the eigenstates of the initial Hamiltonian (which, for our purposes, is Hermitian).
 For a $2\times 2$ system, the outcome can be switched by changing the encircling of the loop from clockwise to anticlockwise.
-The probability of taking one option or the other, which amounts to choosing how the action of the meter takes place, needs to be given by the initial state so as to recover the usual predictions of quantum mechanics.
-This makes the dynamic equation also non-linear, which is consistent with other studies Mertens et al.
+Crucially, to comply with experimental observations, the probability of the interaction selecting a clockwise versus anti-clockwise loop must be imposed according to Born’s rule, determined by the amplitudes of the system’s state vector immediately prior to the interaction.
+As discussed, incorporating this state-dependent probability selection introduces the necessary non-linearity into the description of the measurement process Mertens et al.
 (2021).
-Chiral mode conversion has been extensively studied for $2\times 2$ systems, but there is still much to understand for systems with more degrees of freedom.
-Another mechanism that may come handy is the non-Hermitian skin effect Martinez Alvarez et al.
-(2018); Yao and Wang (2018), which essentially leads to an anomalous localization at one edge of a pristine lattice.
-In some cases one might tune the non-Hermitian term to select a few or even a single eigenstate.
-There has been a surge of interest in this phenomenon, which has been shown to have a topological origin Okuma et al.
-(2020); Zhang et al.
-(2020), see for example Okuma and Sato (2023); Bergholtz et al.
+Conceptually, our analysis utilizing parameter loops around exceptional points offers insight into two key aspects of quantum measurement.
+Firstly, it provides a dynamical mechanism for state selection based on path topology (chirality).
+Secondly, and perhaps more fundamentally, the dynamics inherent to the non-Hermitian evolution during the loop naturally establish the eigenstates of the initial Hamiltonian $\mathcal{H}_{0}$ as the unique stable outcomes, offering a potential mechanistic explanation for their privileged status in this context.
+By signaling how the system is dynamically driven towards these states before one is selected, our framework provides a more complete picture of the collapse process.
+Chiral mode conversion has been extensively studied for $2\times 2$ systems, but there is still much to understand for systems with more degrees of freedom, where other mechanisms such as the non-Hermitian skin effect Martinez Alvarez et al.
+(2018); Yao and Wang (2018) (see also  Bergholtz et al.
 (2021); Lin et al.
-(2023) and references therein.
-We hope that our discussion may trigger a fresh look to this enduring enigma.
-Perhaps the challenge ahead is to take the measurement out of the quantum realm and down to an experiment with classical objects.
-Acknowledgements.
+(2023) and references therein) may come handy.
+As a corollary and perhaps the most immediately useful outcome of our discussion, one can envision the simulation of quantum measurements using classical objects: photonic setups with classical light, electrical circuits, and many others.
+For the case of photonic waveguide arrays Szameit et al.
+(2011); Vicencio et al.
+(2025), the direction of propagation $z$ naturally takes the role of time $t$ Rechtsman et al.
+(2013), owing to the direct analogy between the paraxial wave equation for the light envelope and the time-dependent Schrödinger equation.
+The effective Hamiltonian maps onto the structure’s properties: coupling $J(z)$ via waveguide spacing, detuning $\delta(z)$ via refractive index $n(x,y,z)$, and non-Hermitian terms $\gamma(z)$ via engineered gain/loss (e.g., complex $n$ or coupling to lossy elements).
+Since these parameters can be precisely modulated along $z$ during fabrication (e.g., via laser writing), static waveguide circuits can be created to implement specific time-dependent Hamiltonians, including the parameter loops around exceptional points central to our mechanism.
+This allows for direct observation of the resulting dynamics, including testing the chiral state selection by comparing structures fabricated with opposite loop winding directions along $z$.
+We hope that this may trigger a fresh look at this enduring enigma. Perhaps the challenge ahead is to take the measurement out of the quantum realm and down to an experiment with classical objects.
+V Acknowledgments
+We thank Joaquín Márquez Delgado, Diego Bautista Avilés, Igor Gornyi, Yuval Gefen, and Mikhail Kiselev for useful discussions, and C.
+Cormick, Michael Berry, Franck Laloë and Quancheng Liu for useful comments.
 L.E.F.F.T.
-thanks Joaquín Márquez Delgado, Diego Bautista Avilés, Igor Gornyi, Yuval Gefen, and Mikhail Kiselev for useful discussions, and C.
-Cormick for useful comments.
-L.E.F.F.T.
-acknowledges financial support by ANID FONDECYT (Chile) through grant 1211038, The Abdus Salam International Center for Theoretical Physics and the Simons Foundation, and by the EU Horizon 2020 research and innovation program under the Marie-Sklodowska-Curie Grant Agreement No.
+acknowledges financial support by ANID FONDECYT (Chile) through grants 1211038 and 1250751, The Abdus Salam International Center for Theoretical Physics and the Simons Foundation, and by the EU Horizon 2020 research and innovation program under the Marie-Sklodowska-Curie Grant Agreement No.
 873028 (HYDROTRONICS Project).
@@
 This work is also supported by MICIN with European funds‐NextGenerationEU (PRTR‐C17.I1) and by 2021 SGR 00997, funded by Generalitat de Catalunya.
+VI Author Contributions
+Building on their established research partnership, L.
+E.
+F.
+F.
+T.
+and S.
+R.
+collaborated on this new line of inquiry starting with a question posed by S.
+R..
+For this paper, L.
+E.
+F.
+F.
+T.
+developed the novel concept, led the project, and drafted the manuscript.
+S.
+R.
+contributed valuable insights and critical feedback throughout the process.
+Both authors were involved in refining the ideas and revising the final manuscript.